OpenStudy (anonymous):

According to the Rational Root Theorem ,Which Could be a factor of the polynomial F(x)=60x^4+86x^3-46x^2-43x+8? A)x-6 B)5x-8 C)6x-1 D)8x+5

OpenStudy (anonymous):

someone please help I'm so confused

OpenStudy (triciaal):

one approach if a factor then there is no remainder find the value of x when potential factor = 0 use this value in f(x) if the value computed = 0 then confirmed a factor. another approach synthetic division F(x)=60x^4+86x^3-46x^2-43x+8 + + - - + no yes no yes 2 real positive roots + - -++ yes no yes no 2 real negative roots

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

i think it might be c but im not sure

OpenStudy (triciaal):

if (x-6) is a factor then when x = 6 f(6) = 0 if (6x-1) is a factor then when x = 1/6 is used f(1/6) = 0 etc

OpenStudy (anonymous):

huh sorry im not good at math at all

OpenStudy (triciaal):

this is not using the theorem but I am checking option B 60x^4+86x^3-46x^2-43x+8? 60(8/5)^4 + 86(8/5)^3 - 46(8/5)^2-43(8/5) +8

OpenStudy (triciaal):

I used the graph it confirmed the answer should be B

OpenStudy (triciaal):

the instruction is to use the theorem which is not hard but need to remember how

OpenStudy (triciaal):

the decimal values for x are -1.6, -7.07, 0.707 and 0.167 according to the graph

OpenStudy (triciaal):

the first term is 60x^4 this means will have 4 roots and the product of the coefficients = 60 dividing by x^2 the remainder is -43x + 8 divide by x to get -43 remainder = 8 option B ends in 8

Directrix (directrix):

A statement of the RR Theorem is here; http://www.mathwords.com/r/rational_root_theorem.htm For what it's worth, here's a link which connects to a sample problem to which the Rational Root Theorem has been applied. http://www.purplemath.com/modules/rtnlroot.htm

Directrix (directrix):

C) 6x -1 @17merenfrow0412 The RR theorem offers 1/6 as a possible root. Run it through the given equation by synthetic division to see if it yields 0. As a factor, the root 1/6 would correspond to (6x - 1). 6x -1 = 0 x = 1/6